OpenStudy (anonymous):

Can some one help me. Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

man invn177 you are such a big help. Thank you so much for your time and helping me.

OpenStudy (anonymous):

ok first distribute on both sides... 4x-24 = -2x - 6... then add 6 to both sides... 4x-18=-2x. then subtract 4x from both sides. -18 = -6x... the divide both sides by -6... x=3!

OpenStudy (agreene):

4(x-6)=-2(x-3) distribute 4 and -2 4x-24=-2x+6 add 2x and 24 to both sides 6x=30 divide by 6 x= 5

OpenStudy (anonymous):

invn177 can you help me with one more but tell me if i am doing it right

OpenStudy (anonymous):

sorrsorry i messed up a negative sign agreene is right x = 5

OpenStudy (anonymous):

i wil put it in the question box hold on

OpenStudy (anonymous):

and sure no problem im just bored at work so i get on to do some math, im a nre i know ahha. but ya post it up i got you

OpenStudy (anonymous):

ok]